The inverse demand curve a monopoly faces is p = 100-2Q.

The​ firm's cost curve is C(Q)=30+6Q.

What is the​ profit-maximizing solution?

The​ profit-maximizing quantity is _____. (Round your answer to two decimal​ places.)

The​ profit-maximizing price is $_____ ​(round your answer to two decimal​ places.)
Answer #1

Ans) The​ profit-maximizing quantity is 23.5

The​ profit-maximizing price is $ 53

TR= PQ = 100Q - 2Q^2

MR = dTR /dQ = 100 - 4Q

MC = dTC /dQ = 6

At Profit maximization, MR = MC

100 - 4Q = 6 implies Q = 23.5

P = 100 - 2*23.5 = 53